# Self-Admitted GenAI Usage in Open-Source Software

### v2

@@ 7.2 GenAI Risks and Integration Concerns
Code quality issues have emerged as another significant risk factor closely related to security concerns. Siddiq et al. [40] used the DevGPT dataset to assess the quality of ChatGPT-generated code and found that such code suffers from issues including undefined variables, improper documentation, and security vulnerabilities related to resource management. These quality concerns extend across different programming contexts, as demonstrated by Moratis et al. [41], who analyzed 144 JavaScript code blocks generated by ChatGPT and found that approximately one-quarter of AI-written code blocks contained one or more violations. They observed that approximately 50% of the violations related to best practices, 37% related to code style issues, and 12% were classified as errors-prone violations. Quality concerns increase when considering code modification versus creation. Rabbi et al. [42] analyzed 1,756 AI-generated Python code snippets, systematically distinguishing between code created from scratch and modified code. They found that code modified using ChatGPT more frequently suffers from quality issues compared to ChatGPT-generated code. This pattern suggests that different types of AI assistance may require different governance approaches. Furthermore, Zhang et al. [43] identified code smells in Kubernetes manifest files generated by AI tools, showing that quality concerns extend beyond traditional programming tasks to infrastructure-as-code artifacts.
The successful adoption of GenAI tools requires substantial organizational changes that address both technical and human factors. Sauvola et al. [44] studied the challenge of developer skill adaptation to generative AI, identifying significant skill-gap challenges where developers lack necessary AI expertise. Their findings underline the need for strategic investment in education and training programs to develop new competencies in prompt engineering, AI output validation, and human-AI collaboration techniques. These organizational challenges have also led researchers to investigate GenAI adoption patterns. Russo et al. [45] developed the Human-AI Collaboration and Adaptation Framework, a theoretical model designed to understand and predict GenAI tool adoption in software engineering. They found that compatibility factors—particularly, how well AI tools integrate within existing development workflows—serve as the primary driver of organizational adoption decisions. This finding challenges conventional technology acceptance theories [46], as traditional factors, such as perceived usefulness, social influence, and personal innovativeness, proved less influential than expected in determining GenAI adoption patterns.
@@ 7.3 GenAI Impact on Software Development
Our study adds to the existing body of knowledge by analyzing self-admitted GenAI usage across 250,000+ open-source repositories and conducting a study of code churn pre- and post-GenAI adoption, thus contributing valuable insights on how open-source projects use GenAI tools and how their usage impacts development activity.
@@ 8.1 Construct Validity
When calculating code churn, we used the first explicit mention of GenAI tools as a proxy for adoption timing, which may not accurately reflect when projects actually began using GenAI. However, our generous analysis window of 360 days before and after this point helps accommodate potential discrepancies in adoption dates. Although we focused on only one measure of code quality, the relevance of code churn as a metric was motivated by industry research. The GitClear 2024 report [11] documented a rise in churn from 4.5% in 2023 to 5.7% in 2024, coinciding with the proliferation of GenAI-assisted development. This increase correlates with two related trends: a 39.9% decline in “moved” code (indicating reduced refactoring) and a 17.1% rise in “copy/pasted” code. Previous research links these patterns of less reuse and more duplication to higher defect rates and technical debt [50, 51, 52]. Future work could expand our analysis by considering additional metrics.
@@ 8.2 Internal Validity
The number of policies and guidelines we analyzed and the number of survey responses we received was relatively low. However, even this limited data revealed diverse regulation approaches and opinions, motivating future research.
@@ 8.3 External Validity
We restricted our analysis to public repositories hosted on GitHub, focusing on five popular programming languages. Our sample of repositories might not represent GenAI usage patterns in other repositories, programming languages, or industrial software projects. However, the selected languages represent the most commonly used languages according to the 2024 GitHub Octoverse report [13]. Furthermore, our filtering criteria for engineered software projects ensured that our findings reflect practices in actively maintained software projects. Finally, our focus on ChatGPT and GitHub Copilot might not capture usage patterns of other GenAI tools.
@@ 9 Conclusion
This study introduced self-admitted GenAI usage—explicit references to LLM-based tools such as ChatGPT and GitHub Copilot—as a novel lens for examining how generative AI is used in open-source software development. In our mixed-methods study design, we first mined more than 250,000 GitHub repositories, isolating 1,292 true-positive GenAI mentions across 156 projects. Qualitative open coding of these instances and subsequent card sorting yielded taxonomies of 32 assisted tasks, 10 generated content types, and 11 usage purposes. We complemented this content analysis with a survey of project contributors and a systematic review of 13 project-level policies and guidelines. In addition, we performed a regression-discontinuity analysis of code churn in 149 repositories that contained sufficient data to study the impact of GenAI adoption on open source software projects. We found that:
RQ1 (usage in practice): Developers most frequently use GenAI tools for code generation, natural-language translation, and refactoring; acknowledgment is a common purpose; human follow-up actions underscore the importance of human oversight.
RQ2 (GenAI governance): Project responses range from outright bans on AI-generated contributions to practical acceptance, subject to disclosure or additional review. Developers consider ethical and legal uncertainties around copyright, licensing, and data privacy. Some project guidelines actively encourage GenAI usage for supporting tasks such as pull-request descriptions or documentation.
RQ3 (impact on activity): Contrary to a popular industry report, our repository-specific analysis detects no systematic post-GenAI adoption rise in code churn. Average file- and line-based churn declined modestly across the full commit set, with only a minority of projects exhibiting a significant upward trend.
These results have several implications for software developers. First, the explicit self-admissions we found are a first step towards more transparent AI use. However, the sparsity of contextual metadata (e.g., prompts or model versions) suggests that community conventions for provenance reporting remain to be formed. Second, heterogeneous governance approaches signal that one-size-fits-all policy prescriptions are unlikely to succeed. Instead, project maintainers must calibrate the guidelines and policies to their specific project context. Third, our churn analysis cautions against broad analyses that do not factor in when GenAI adoption in particular projects happened and if and how they regulate GenAI usage. This reinforces the need for empirical project-level monitoring.
For researchers, our curated dataset of annotated GenAI mentions and accompanying taxonomy creates a foundation for automated detectors and studies that mine GenAI mentions on a larger scale. Future work should triangulate churn with complementary metrics (defect density, clone rates, review latency) and expand the analytical lens beyond ChatGPT and GitHub Copilot.
@@ ACKNOWLEDGMENTS
We thank all participants who took the time to complete our survey, providing valuable insights for our research. The research contribution of Fabio Calefato was partially supported by the European Union - NextGenerationEU through the Italian Ministry of University and Research, Projects PRIN 2022 (“QualAI: Continuous Quality Improvement of AI-based Systems”, grant n. 2022B3BP5S, CUP: H53D23003510006).

### v3

@@ 7.2 GenAI Risks and Integration Concerns
Code quality issues have emerged as another significant risk factor closely related to security concerns. Siddiq et al. [39] used the DevGPT dataset to assess the quality of ChatGPT-generated code and found that such code suffers from issues including undefined variables, improper documentation, and security vulnerabilities related to resource management. These quality concerns extend across different programming contexts, as demonstrated by Moratis et al. [24], who analyzed 144 JavaScript code blocks generated by ChatGPT and found that approximately one-quarter of AI-written code blocks contained one or more violations. They observed that approximately 50% of the violations related to best practices, 37% related to code style issues, and 12% were classified as error-prone violations. Quality concerns increase when considering code modification versus creation. Rabbi et al. [31] analyzed 1,756 AI-generated Python code snippets, systematically distinguishing between code created from scratch and modified code. They found that code modified using ChatGPT more frequently suffers from quality issues compared to ChatGPT-generated code. This pattern suggests that different types of AI assistance may require different governance approaches. Furthermore, Zhang et al. [55] identified code smells in Kubernetes manifest files generated by AI tools, showing that quality concerns extend beyond traditional programming tasks to infrastructure-as-code artifacts.
The successful adoption of GenAI tools requires substantial organizational changes that address both technical and human factors. Sauvola et al. [38] studied the challenge of developer skill adaptation to generative AI, identifying significant skill-gap challenges where developers lack necessary AI expertise. Their findings underline the need for strategic investment in education and training programs to develop new competencies in prompt engineering, AI output validation, and human-AI collaboration. These organizational challenges have also led researchers to investigate GenAI adoption patterns. Russo et al. [35] developed the Human-AI Collaboration and Adaptation Framework, a theoretical model designed to understand and predict GenAI tool adoption in software engineering. They found that compatibility factors—particularly, how well AI tools integrate within existing development workflows—serve as the primary driver of organizational adoption decisions. This finding challenges conventional technology acceptance theories [22], as traditional factors, such as perceived usefulness, social influence, and personal innovativeness, proved less influential than expected in determining GenAI adoption patterns.
@@ 7.3 GenAI Impact on Software Development
Our study adds to the existing body of knowledge by analyzing self-admitted GenAI usage across 250,000+ open-source repositories and conducting a longitudinal study of code churn, thus contributing valuable insights on how open-source projects use GenAI tools and how their usage impacts development activity.
@@ 8.1 Construct Validity
The focus on self-admitted usage has implications for our answers to the research questions. Developers may be more likely to acknowledge GenAI usage for mundane tasks such as translation rather than for core development tasks such as implementing complex business logic. Therefore, the observed frequencies might reflect what developers feel comfortable disclosing rather than their actual usage (RQ1). Strict usage policies might lead to fewer self-admitted usages, as contributors might avoid disclosure (RQ2). Finally, unacknowledged usage before the first explicit mention could distort the pre-adoption baseline for our code churn analysis (RQ3). However, at the same time, we consider self-admitted GenAI usage a useful analytical lens for studying the impact of GenAI adoption in open-source software projects. Given our manual validation, the precision of the identified GenAI usages is high, even though the overall recall of all GenAI usages is low.
When calculating code churn, we used the first explicit mention of GenAI tools as a proxy for adoption timing, which may not accurately reflect when projects actually began using GenAI. However, our generous analysis window of 360 days before and after this point and robustness checks (e.g., placebo tests) help accommodate potential discrepancies in adoption dates. Although we focused on only one measure of code quality, the relevance of code churn as a metric was motivated by industry research. The GitClear 2024 report [10] documented a rise in churn from 4.5% in 2023 to 5.7% in 2024, coinciding with the proliferation of GenAI-assisted development. This increase correlates with two related trends: a 39.9% decline in “moved” code (indicating reduced refactoring) and a 17.1% rise in “copy/pasted” code. Previous research links these patterns of less reuse and more duplication to higher defect rates and technical debt [23, 19, 8]. Future work could expand our analysis by considering additional metrics.
@@ 8.2 Internal Validity
The number of policies and guidelines we analyzed and the number of survey responses we received were relatively low. However, even this limited data revealed diverse regulation approaches and opinions, motivating future research.
@@ 8.3 External Validity
We restricted our analysis to public repositories hosted on GitHub, focusing on five popular programming languages. The self-admitted GenAI usage we studied might not reflect general GenAI usage—self-admitted or hidden—in other repositories, programming languages, or industrial software projects. However, the selected languages represent the most commonly used languages according to the 2024 GitHub Octoverse report [12]. Furthermore, our filtering criteria for engineered software projects ensured that our findings reflect practices in actively maintained software projects. Moreover, our results may not generalize to other open-source platforms such as GitLab, which may have different norms and adoption patterns.
The developer survey that we conducted as part of RQ2 received eight responses by contributors of one of the 12 projects that we identified to have explicit GenAI usage policies and guidelines. Their responses might not reflect the views of a broader developer population. Future work should extend this analysis, for example, by analyzing GitHub Discussion threads on AI regulation. Finally, our focus on ChatGPT and GitHub Copilot might not capture the usage patterns of other GenAI tools that were released more recently, e.g., Claude Code.
@@ 9 Conclusion
This study introduced self-admitted GenAI usage—explicit references to LLM-based tools such as ChatGPT and GitHub Copilot—as a novel lens for examining how generative AI is used in open-source software development. In our mixed-methods study design, we first mined more than 250,000 GitHub repositories, isolating 1,292 true-positive GenAI mentions across 156 projects. Qualitative open coding of these instances and subsequent card sorting yielded taxonomies of 32 assisted tasks, 10 generated content types, and 11 usage purposes. We complemented this content analysis with a survey of project contributors and a systematic review of 13 project-level policies and guidelines. In addition, we performed a regression-discontinuity (RDD) analysis of code churn in the 149 repositories that contained sufficient data to study the impact of GenAI adoption on open source software projects. Based on our findings, we derive several actionable implications.
Implications for software developers. The RQ1 results show that developers most often use GenAI for code generation, natural-language translation, and refactoring—with explicit acknowledgment as the dominant purpose. We also observed recurring follow-up actions, including bug fixes, refactorings, reversions, and deletions, triggered by earlier GenAI-generated content. Together with our RQ3 task-level analysis, which shows that generation tasks are more rework-prone than optimization or maintenance, this suggests that GenAI output has a provisional nature. Developers should therefore plan explicit validation and revision steps after GenAI-assisted generation, particularly for code. While acknowledgment of GenAI usage is common (RQ1), contextual metadata such as prompts, model versions, or scope of generation is rarely recorded. This gap points to a concrete opportunity for developers to adopt provenance conventions (e.g., structured comments or commit tags) that better support the corrective practices already observed in the data. The RQ2 findings show substantial variation in project-level GenAI governance. Developers contributing to different projects should therefore avoid assuming uniform norms and instead adapt their GenAI usage and disclosure practices to project-specific guidelines.
Implications for project maintainers. RQ2 shows that maintainers are already actively regulating GenAI usage, but through heterogeneous approaches ranging from bans to selective encouragement. This diversity suggests that generic policies are unlikely to work. Instead, maintainers should align GenAI guidelines with project-specific factors such as contribution patterns, review capacity, and risk tolerance. The RQ3 results show no systematic increase in code churn after GenAI adoption, contradicting prominent industry claims. While some repositories exhibit increased churn, many show stable or decreasing trends, and effects vary widely across projects. This suggests that restrictive policies motivated solely by assumed quality degradation are not well supported by the evidence. Project-level monitoring, e.g., tracking churn or related indicators before and after GenAI adoption, offers a project-specific and evidence-based alternative. The findings for RQ1 highlight that GenAI is frequently used for PR descriptions and documentation, which are comparatively low-risk artifacts. Maintainers can act on this by explicitly encouraging GenAI usage in these areas to improve communication and review efficiency while limiting exposure to higher rework costs.
Implications for tool builders and platform providers. RQ1 shows that developers already engage in voluntary disclosure of GenAI usage when the cost is low, indicating that the main limitation to transparency lies in the absence of supporting mechanisms rather than in developer opposition. Tool builders can act on this by offering built-in, optional disclosure mechanisms such as automatic annotations or commit templates that integrate with existing workflows. RQ3 indicates that GenAI-assisted code generation is more rework-prone than other uses. Tools could respond by explicitly supporting post-generation validation, for example, through prompts or workflow affordances that encourage human review of generated code. At the platform level, support for project-specific GenAI policies, such as configurable disclosure requirements in pull requests, could help operationalize the heterogeneous governance approaches observed in RQ2.
Implications for researchers. Our curated dataset of 1,292 self-admitted GenAI mentions and the taxonomy derived in RQ1 provide a foundation for scaling empirical studies via automated detection and large-scale mining. Researchers can build on this work by developing detectors for self-admitted GenAI usage and using them to study adoption and disclosure practices at scale. RQ3 underscores the importance of methodological granularity. Analyses that ignore project-specific adoption points or governance contexts risk drawing misleading conclusions. Future studies should therefore favor repository- and task-level designs when assessing the impact of GenAI usage, rather than relying on aggregate trends. Because this study focuses on version-controlled artifacts, future research should extend the analysis to adjacent artifacts such as pull-request discussions and review comments. Such extensions would naturally build on our dataset, taxonomy, and churn analysis, helping to complete the picture of how self-admitted GenAI usage shapes collaborative software development.
@@ ACKNOWLEDGMENTS
We thank all participants who took the time to complete our survey, providing valuable insights for our research. The research contribution of Fabio Calefato was partially supported by the European Union - NextGenerationEU through the Italian Ministry of University and Research, Projects PRIN 2022 (“QualAI: Continuous Quality Improvement of AI-based Systems”, grant n. 2022B3BP5S, CUP: H53D23003510006). Tao Xiao is supported in part by JSPS Grant-in-Aid for JSPS Fellows 23KJ1589 and the Kayamori Foundation of Informational Science Advancement.